User Tools

Site Tools


wibed:doc_eloi:project_log

Project LOG

This project started at 15th of January but the transcription of the work done in this log started at 27th of January. However, all the work done is in this wiki in the corresponding sections.

27th January - 2nd February:

  • Updating the nodes to current version of WIBED firmware (using last commits in the repository)
  • Trying to deploy some nodes in E classrooms in A5/A6 Building (in the dropped ceiling)
  • Contacting the Libremap mail list in order to discuss how to implement the elevation in the system
    • The main developers says that it is too much work to do in order to use 3D visualization in Libremap and that it could be a complete Master/PhD project.
    • I asked them that maybe a easy plugin that paints the node in different colours according to its location's height. (waiting for a response)
  • Maintenance managers are away and there is no possibility to discuss how to deploy the nodes in the dropped ceiling.

3rd February - 9th February:

  • Maintenance manager meet me and we have been looking in the both E classroms of the A6 building. Finally I have placed a node (a5fa34) in a maintenance bathroom in its dropped ceiling (set in a temporary place)
    • Maintenance manager says that dropped ceilings in the classrooms may be too thin to fit the nodes and its radios. We agreed to take a proper look to this in the future
  • Internet is down and there is no possibility to check the correct behaviour of the mesh.
  • Internet connection has been restored and the new nodes can be seen but, because of the reinforced cement, the signal is too poor and there are disconnections constantly (we need to check if adding 1 more node in the A6-E01 or near there, we fix that).
  • 10 new Radios (USB dongles) and 10 16GB USB had arrived. We need one more USB Dongle and there are 4 16GB USB spares.
  • We are testing an update for the wibed firmware (using the last updates of the OpenWRT distribution instead of an old stable version). Seems that in the node 3e9e8e works fine but we will continue testing.
  • If all tests are OK we will add a new branch into the repository with this version (unstable?)
    • This branch (merge_trunk) has been added to the repository
  • We are seeing too much problems with the Radio2 being the BATMAN mgmt main device.
    • The gateway falls because of the Radio2 constant freezes
    • Some Radio2 of other nodes freezes too.
    • Manos said that maybe it would be better to use the Radio0 as the BATMAN management device and the Radio2 to deploy the experiments.
      • Deploying some experiments, the radio2 seems to work perfectly as the experiment radio and the radio0 have no problems with the management. We have to check it better but seems that next commint&image will use these changes.
      • I am configuring an image (in local) with the changes and will be flashed in some devices tomorrow to test how they work.

10th February - 9th February:

  • We have changed some nodes to the new configuration + firmware and seems to be working fine.
  • The server has been upgraded to the new Manos version and the experiments and behaviour seems correct.
    • We are currently testing the experiments and the state of the node once the experiment is finished (One wibed developer says that he find old experiment files in the overlay) at this moment we can not reproduce the issue and all seems correct.
    • The node e4b63a had some kind of error maybe related with the USB overlay. After a lot of checks, sysupgrades and experiments it has been fixed using this:
      • 1st command: firstboot
      • 2nd command: reboot -f
  • We are facing problems to deploy the experiments: the nodes are not working properly with the wibed different status. Some nodes changes to RESETING or ERROR and each experiment needs some work on each node facing problems. Most common solution is to try with:
    • uci set wibed.general.status=1
    • uci set wibed.experiment.exp_id=0
    • wibed-reset
  • or just try if this works:
    • wibed-config
  • We have tried to hide the SSIDs of adhoc management wifis. To hide the SSID is not compatible with the adhoc mode of the nodes.

17th February - 23th February

  • Manos have told to deploy a mini-testbed in the C6 in order to do an experiment with 10-12 nodes. This testbed may be huge and it will compare the behaviour of some networking protocols like 802.11s, Batman-adv or others.
  • We have deployed some new nodes: 2 of them are deployed in the A6 building on the E floor. They has been setted in a ventilation space (outdoors) in both sides of the building. One more node is deployed in the A5 floor -1 in a maintenance hidden wardrobe near the GW.
  • The GW has been reflashed to qmp. The Lab GW has been changed to another node and the A5 GW has been recovered and updated to latest version.

25th February - 2nd March

  • We have deployed a node in the A5 -1 floor in the hidden wardrobe above the stairs → 6041
  • We have deployed 2 nodes in the A5 1st floor in the wardrobes in front of the bathrooms (one of each side) → 99c4 and 0cee
    • We faced problems because maintenance workers connected the wall socket to a timed plug (lights)
  • Someone unplugged the node 3e9e7c. Now is working again. Now we have connectivity from A5 -1 floor to A5 1st floor (deployment of a new node in the E floor expected to be done next week).

3rd March - 9th March && 10th March - 16th March

  • The deployment of the node in the A5-E02 floor (in the ventilation compartment) is stopped because there is an extraction system that works in summer and it will damage our node. The deployment is still being studied.
  • Wifix experiments started in the A5-A6 deployment. There are some connectivity problems that must be solved.
  • 4 new nodes will be deployed soon (12~14th March) that will bring A5 and A6 2nd floor connectivity. Is expected that this nodes will have good connectivity with 1st floor nodes and be able to deploy experiments without problems.

17th March - 23rd March

  • The nodes {163e8e,1642be,f41a44,faaf86} are deployed and working. There is still an issue with the node 163e8e because it has much more ping that expected. We are studying why this happens (noise/signal/reinforced cement…} and how to fix it.
  • We are studying to deploy a TP-Link WDR4900 (N900) router with a powerful antennas in the E-floor in order to improve connectivity.
  • We are working on the next update of the firmware of the node. That will improve antenna tx-power capacity and update the batman-adv protocol version.
  • Still studying how to improve the connectivity in the most conflictive areas.

24th March - 30th March

  • Testing the first implementations of the next version of the firmware (all works correctly at the moment). Trying to fix some bugs of the old version that are not working properly.
  • Improving the geolocation configuration of the nodes (adding functionalities to the server and firmware).
  • The new last_trunk firmware version is released and available in the repository.
  • The Wibed team is working in the experiments that will be used in order to validate the platform and focusing in the most urgent tasks.
  • This week we have been focused also in a bug that do not allow the routers to set its country, so the antenna power is not properly set. The problem is related with the ath9k_htc (USB Atheros Radio) driver. We are studying how to fix this problem.

31st March - 6th April

  • Checking how to improve the TXPower and fix the problem with the Ath9k-htc driver (the usb radio driver). (Recurrent problem of previous weeks)
  • Improved the geolocation configuration and settings. Now this can be set in the server and automatically updated in the node (currently in beta, next week in the stable server). Added some scripts in order to improve this behaviour.
  • Improving the txpower configuration.

7th April - 13th April

  • Started a new branch in the routing git repository in order to improve the configuration of the nodes (last batman-adv 2014.1.0 has some bugs that stubs the WiFi interfaces when some configurations are changed).
  • Updated all the nodes in the A5/6 buildings in order to have all the deployment in the latest wibed version.

14th April - 20th April

  • All buildings and services of the university are closed. We have performed some tasks but we could not access to the nodes.
  • Improved some capabilities of the wibed packages.
  • Planning and studying the expansion of the network.

21st April - 27th April

  • Discussing the proposal of network expansion.
  • Discussing to add new type of routers (Alix).
  • Resolving some issues with the nodes (University services and buildings are opened again).

28th April - 4th May

  • Working on the new deployment.
  • Working on the Wireless Battle of the Mesh V7 in Leipzig.
  • Working on the experiments that will validate wibed as a trusty and reliable platform.

5th May - 11th May

  • Working on the new deployment.
  • Working on the Wireless Battle of the Mesh V7 in Leipzig.
  • Working on the experiments that will validate wibed as a trusty and reliable platform.
wibed/doc_eloi/project_log.txt · Last modified: 2014/05/06 12:13 by eloicaso